OpenStudy (anonymous):

Using the two functions listed below, insert numbers in place of letters A,b,c and d so that f(x) and g(x) are inverse

OpenStudy (anonymous):

what are the two functions...?

OpenStudy (anonymous):

stull need help?

OpenStudy (anonymous):

Here is the full question Part 1. Using the two functions listed below, insert numbers in place of the letters a, b, c, and d so that f(x) and g(x) are inverses. f(x)= x+a b g(x)=cx−d Part 2. Show your work to prove that the inverse of f(x) is g(x). Part 3. Show your work to evaluate g(f(x)). Part 4. Graph your two functions on a coordinate plane. Include a table of values for each function. Include 5 values for each function. Graph the line y = x on the same graph.

OpenStudy (anonymous):

@BlooooooBlahBLIIIIK

OpenStudy (anonymous):

@rheaanderson341

OpenStudy (jdoe0001):

hmmm have you covered inverse functions yet?

OpenStudy (anonymous):

Yes @jdoe0001

OpenStudy (jdoe0001):

hmmm need to dash in a bit but hmm what I'd do is, pick two random values for "a" and "b" say 2 and 3 so f(x) = x+ab f(x) = x+(2 * 3) f(x) = x +6

OpenStudy (jdoe0001):

so, if we use 2 and 3 for "a" and "b" we end up with a function in "x" terms so to find f(x), or "y" we can simply use some "x" value, say 4 f(x) = x +6 x=4 f(x) = 4+6 f(4) = 10

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

It's f(x)=x+a/b

OpenStudy (jdoe0001):

so... all that is thus far f(x) only BUT recall that the domain of one function, is the range of the inverse function that is if we use x= 4, as INPUT, and our OUTPUT is 10 then the inverse would have to take in, 10 and output 4

OpenStudy (anonymous):

So y=x+2/3?

OpenStudy (jdoe0001):

yes... using a/b then get a random value for "x", to get some "y"

OpenStudy (anonymous):

Okay so would that be the inverse function for fx?

OpenStudy (jdoe0001):

so you get some OUTPUT from some INPUT and reverse that on g(x) IF g(x) is to be the inverse function, the OUTPUT of f(x), will serve as the INPUT for g(x)

OpenStudy (jdoe0001):

and if you use those INPUT and OUTPUT values from f(x) into g(x) you can then solve for either "c" or "d" and use a random "d", to get a "c", or the other way around and g(x) will end up being the inverse because of the INPUT and OUTPUT swap

OpenStudy (jdoe0001):

anyhow... dashing :)

OpenStudy (anonymous):

Okay.... So how would I get the equations